OpenStudy (anonymous):

Last question someone help? or @dabest21

OpenStudy (anonymous):

Which equation represents the relationship shown in the table?

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

choices @AnasAlazzawi

OpenStudy (anonymous):

Are there choices @AnasAlazzawi

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

1) y equals negative one-half x minus one 2) y equals x plus one 3) y equals one-half x minus one 4) y equals one-half x plus one

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

1. Y= -1/2 x-1 2. Y= x+1 3. y= 1/2x-1 4. y=1/2x+1

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

its not y=x+1 because (0,-1) would be 0,1

OpenStudy (anonymous):

that's in the answer choices of the quiz

OpenStudy (anonymous):

So is it 1,2 or 4

OpenStudy (anonymous):

just pick a coordinate and plug its x in for x and y in for y in each equation ill do it with u

OpenStudy (anonymous):

lets try (-4,-3)

OpenStudy (anonymous):

for 1 it would be -1/2*-4-1 2-1 1 does -3=1

OpenStudy (anonymous):

no so u try equation 3 on your own and ill check its y=1/2*x-1 so substitute -4 for x and try it y=1/2*-4-1 so u do that @anas

OpenStudy (anonymous):

imean @AnasAlazzawi

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So which one is it

OpenStudy (anonymous):

I'm highly confused

OpenStudy (anonymous):

did u do 1/2*-4-1

OpenStudy (anonymous):

No

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

its-3

OpenStudy (anonymous):

yes so that proves number 3 is the answer cause the point was -4,-3 and -3=-3

OpenStudy (anonymous):

Don"t forget to medal and fan if you already haven't. Plus mention me for more help @AnasAlazzawi